# Multi-Rank Subspace Change-Point Detection for Monitoring Robotic Swarms

### v1

@@ 5 Numerical experiments
5.1 Simulations
In this section, we evaluate the Subspace-CUSUM procedure on synthetic data. We simulate sequences of $k$-dimensional Gaussian observations $x_{t}\sim N(0,\Sigma_{t})$, where $\Sigma_{t}=I_{k}$ for $t\leq\tau$ before an unknown changepoint $\tau$ and $\Sigma_{t}=\Sigma_{1}$ for $t>\tau$. We consider several spiked covariance scenarios for $\Sigma_{1}$:
  • Rank-1 spike: $\Sigma_{1}=I_{k}+\lambda uu^{T}$, where $u$ is a unit-norm eigenvector. We consider both a dense spike where $u=\frac{1}{\sqrt{k}}(1,1,\dots,1)^{T}$ and a sparse spike where $u=e_{1}$.
  • Rank-$d$ spike: $\Sigma_{1}=I_{k}+U\Lambda U^{T}$ with a $d$-dimensional orthonormal subspace $U=[u_{1},\dots,u_{d}]$. We consider uniform spikes $\Lambda=\lambda I_{d}$ where all $d$ signal eigenvalues are equal, and non-uniform spikes $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$ with eigenvalues $\lambda_{1}>\lambda_{2}>\cdots>\lambda_{d}>0$ of varying strength.
In our experiments, we benchmark Subspace–CUSUM against the oracle Exact CUSUM, which has full knowledge of the post-change subspace, and three practical baselines: Eigenvalue Shewhart Chart (xie2020sequential, ), Hotelling’s $T^{2}$ chart (hotelling1947multivariate, ), and GLR (Generalized Likelihood Ratio) procedure.
Table 1 demonstrates the thresholds $b$ for the Subspace-CUSUM procedure that yield an ARL of approximately 5000 under the pre-change distribution, calibrated by Monte-Carlo simulation. We show results for several combinations of ambient dimension $k$, subspace rank $d$, and sliding window length $w$. In all cases, the post-change signal strength was set to $\Lambda=I_{d}$ with a lower bound of componentwise SNR of $\rho_{\min}=0.5$.
Table 1: Threshold $b$ for stopping rules corresponding to ARL = 5000, calibrated by Monte-Carlo simulation: $\Lambda=I_{d}$, $\rho_{\min}=0.5$.
$w$ $20$ $50$ $100$
@@ 5.1 Simulations
(a) $\sigma^{2}=2$
(b) $\sigma^{2}=1$
(c) $\sigma^{2}=0.5$
[…]
@@
Figure 1 illustrates the trade-off between ARL and EDD for three online detectors under the uniform spike model with $k=10$, $d=2$, and window length $w=50$. Each subplot corresponds to a different noise variance of $\sigma^{2}=2,1,0.5$. Across all noise levels, the Exact CUSUM attains the smallest expected detection delay (EDD) for any prescribed average run length (ARL), corroborating its theoretical optimality. When the signal-to-noise ratio is low (high noise), Subspace–CUSUM closely tracks this oracle performance, suffering only a modest additional delay. By contrast, the Eigenvalue Shewhart chart incurs substantially larger delays under weak signals, since its one-shot decision statistic reacts slowly to small shifts. As the signal-to-noise ratio increases (low noise), the Shewhart-type chart begins to outperform Subspace–CUSUM, reflecting the well-known fact that Shewhart charts excel at detecting large, abrupt changes (montgomery2020introduction, ).
Table 2: Simulated detection delays corresponding to ARL calibrated to 5000: $\Lambda=I_{d}$. CUSUM, S-CUSUM and EigChart stand for the Exact CUSUM, Subspace CUSUM and Eigenvalue Shewhart Chart, respectively. Hotelling’s $T^{2}$ and GLR are omitted due to EDD $\approx 5000$.
@@
CUSUM S-CUSUM EigChart CUSUM S-CUSUM EigChart CUSUM S-CUSUM EigChart
5 2 52.9 159.6 850.2 20.1 77.1 90.6 8.4 63.7 40.1
5 3 37.9 122.3 722.8 14.4 76.2 85.8 6.0 57.8 42.9
10 2 52.5 196.3 1225.7 20.2 86.8 114.1 8.4 59.9 45.2
10 3 38.0 153.8 1043.9 15.8 75.0 100.2 6.0 62.7 50.3
20 2 52.4 349.1 2072.5 20.1 106.9 185.6 8.4 63.2 47.1
20 3 38.1 247.5 1802.0 14.4 101.2 166.4 6.0 63.2 53.7
@@
Table 2 reports the numerical values of the EDD for each detector when its threshold is tuned to yield an ARL of 5000 under the pre-change model. We consider three ambient dimensions $k\in\{5,10,20\}$ and two subspace ranks $d\in\{2,3\}$, across the same three noise variances as in Figure 1. In every configuration, the Exact CUSUM attains the smallest delay, matching its oracle status. Subspace–CUSUM delivers a delay that is only slightly larger, with the gap narrowing as the signal-to-noise ratio improves (i.e. $\sigma^{2}$ decreases). The Largest-Eigenvalue Shewhart chart again shows substantially higher delays at low SNR, though it becomes more competitive for stronger signals. These results corroborate the trade-offs between ARL and EDD seen in Figure 1 and demonstrate that Subspace–CUSUM offers a great balance between false-alarm control and rapid detection in multi-rank spiked covariance scenarios. Note that we have omitted Hotelling’s $T^{2}$ and the GLR procedure from Table 2, as both methods yielded enormous EDDs that closely matched the target ARL of 5000 across almost all configurations.
5.2 Real data
5.2.1 Swarm behavior data
Figure 2: Detection statistics (scaled to 0-1) over time for the swarm dataset, as produced by the different change-point detection methods.
@@ 5.2.1 Swarm behavior data
We consider a swarm trajectory dataset consisting of $T=10{,}001$ sequential frames of collective behavior. In each frame $t=1,\dots,T$, there are $n=100$ agents moving in a 2D plane, and we record each agent’s two-dimensional positions $P_{t}=\{p_{t1},\ldots,p_{tn}\}$ and velocity vectors $V_{t}=\{v_{t1},\ldots,v_{tn}\}$, where $p_{ti},v_{ti}\in\mathbb{R}^{2}$, $i=1,\ldots,n$. We can construct an affinity graph, where each agent corresponds to a node, and we can observe a time series of nodal features $\hat{X}_{t}=[P_{t};V_{t}]\in\mathbb{R}^{4\times n}$. Let $\tilde{X}_{t}$ be the position-centered and normalized nodal feature values from $\hat{X}_{t}$, which is used as the input for change-point detection.
@@
We compare Subspace-CUSUM with three other baseline approaches, including GLR based on vectorized data, Eigenvalue Shewhart Chart, and Subspace Tracking (zhang2023spectral, ). Figure 2 illustrates the result. Subspace-CUSUM yields some clearly identifiable change points, as evidenced by a sharp rise in its detection statistic. Although no ground-truth change time is available for this real dataset, the timing of the detected change-point closely matches Subspace Tracking and an independent result obtained via the iso-mirror method of chen2023discovering . Notably, Subspace-CUSUM and chen2023discovering are founded on entirely different statistical and algorithmic principles. The concordance between detected change-points strongly suggests that both methods are capturing the same underlying network transition.
@@ 6 Conclusion
We present the multi-rank Subspace-CUSUM algorithm for sequential detection of covariance structure changes in high-dimensional streams. This method extends previous subspace monitoring approaches to handle multiple concurrent spikes, offering more powerful detection capabilities in scenarios involving multiple simultaneous low-rank signals. We provide a theoretical foundation that guides the selection of key parameters to control the average run length, and we show through analysis and simulations that the proposed detector can reliably achieve the targeted performance.
Looking ahead, a key open problem is establishing the theoretical optimality of the proposed procedure. While our design is guided by asymptotic heuristics and shows strong empirical performance, a formal proof of optimality remains to be shown. Future work may also extend our framework to more general settings, such as non-Gaussian noise or time-varying covariances, and explore adaptive strategies for selecting the subspace dimension.
@@ Acknowledgement
This work is supported by the Office of Naval Research [N000142412278].
@@ Appendix A Proofs
For notational simplicity, we here denote by $\hat{u}_{i}$ the $i$-th leading unit-norm eigenvector of the sample covariance $\hat{\Sigma}_{t+w}$, which is the $i$th column of $\hat{U}_{t+w}$. Under the spiked covariance model with distinct signal strengths $\lambda_{i}$, we have the following asymptotic normality of $\hat{u}_{i}$ [1, 17]:
$\sqrt{w}(\hat{u}_{i}-u_{i})\stackrel{{\scriptstyle d}}{{\to}}N(0,\Xi_{i}),$
where the $k\times k$ covariance matrix $\Xi_{i}$ admits the spectral decomposition
@@
This $\Xi_{i}$ can be amounted to
$\Xi_{i}=\frac{1+\rho_{i}}{\rho_{i}^{2}}(I_{k}-UU^{T})+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{d}\frac{(1+\rho_{i})(1+\rho_{j})}{(\rho_{i}-\rho_{j})^{2}}u_{j}u_{j}^{T}.$
Using this result, we have as $w\to\infty$,

### v2

@@ 6 Numerical experiments
We evaluate the empirical performance of the MRS-C procedure based on both synthetic and real datasets in order to assess false-alarm control, detection delay, and robustness of the parallel process.
6.1 Simulation study: ARL calibration and baseline comparison
We begin with synthetic experiments that assess ARL calibration and compare detection delays across competing methods under controlled parameters. We simulate sequences of $k$-dimensional Gaussian observations $x_{t}\sim N(0,\Sigma_{t})$, where $\Sigma_{t}=\sigma^{2}I_{k}$ for $t\leq\tau$ before an unknown change-point $\tau$ and $\Sigma_{t}=\Sigma_{1}$ for $t>\tau$. We consider several spiked covariance scenarios for $\Sigma_{1}$:
  • Rank-1 spike: $\Sigma_{1}=\sigma^{2}I_{k}+\lambda uu^{T}$, where $u$ is a unit-norm eigenvector. We consider both a dense spike where $u=(1,1,\dots,1)^{T}/\sqrt{k}$ and a sparse spike where $u=e_{1}=(1,0,\dots,0)^{T}$.
  • Rank-$d$ spike: $\Sigma_{1}=\sigma^{2}I_{k}+U\Lambda U^{T}$ with a $d$-dimensional orthonormal subspace $U=[u_{1},\dots,u_{d}]$. We consider uniform spikes $\Lambda=\lambda I_{d}$ where all $d$ signal eigenvalues are equal, and non-uniform spikes $\Lambda=\operatorname{diag}(\lambda_{1},\dots,\lambda_{d})$ with eigenvalues $\lambda_{1}>\lambda_{2}>\cdots>\lambda_{d}>0$ of varying strength.
In our experiments, we benchmark the MRS-C against the oracle Exact CUSUM, which has full knowledge of the post-change subspace, and the Largest-Eigenvalue Shewhart Chart (xie2020sequential).
Table 1 demonstrates the thresholds $b$ for the MRS-C procedure that yield an ARL of approximately 5000 under the pre-change distribution, calibrated by Monte-Carlo simulation. We show results for several combinations of ambient dimension $k$, subspace rank $d$, and sliding window length $w$. In all cases, the post-change signal strength was set to $\Lambda=I_{d}$ with a lower bound of componentwise SNR of $\rho_{\min}=0.5$.
Table 1: Threshold $b$ for stopping rules corresponding to ARL of 5000, calibrated by Monte-Carlo simulation: $\Lambda=I_{d}$, $\rho_{\min}=0.5$.
$w$ 2020 $50$ $100$
@@ 6.1 Simulation study: ARL calibration and baseline comparison
(a) $\sigma^{2}=2$.
(b) $\sigma^{2}=1$.
(c) $\sigma^{2}=0.5$.
@@
Figure 1 illustrates the trade-off between ARL and EDD for three online detectors under the uniform spike model with $k=10$, $d=2$, and window length $w=50$. Each subplot corresponds to a different noise variance of $\sigma^{2}=2,1,0.5$. Across all noise levels, the Exact CUSUM attains the smallest detection delay for any prescribed ARL, corroborating its theoretical optimality. When the signal-to-noise ratio is low (high noise), MRS-C closely tracks the oracle performance, incurring only a modest additional delay. By contrast, the Eigenvalue Shewhart chart incurs substantially larger delays under weak signals, since its one-shot decision statistic reacts slowly to small shifts. As the signal-to-noise ratio increases (low noise), the Shewhart-type chart begins to outperform MRS-C, reflecting the well-known fact that Shewhart charts excel at detecting large, abrupt changes (montgomery2020introduction).
Table 2: Simulated detection delays with ARL calibrated to 5000 under $\Lambda=I_{d}$. “CUSUM” and “LESC” denote the Exact CUSUM and the Largest-Eigenvalue Shewhart Charts, respectively (standard errors in parentheses).
@@
CUSUM MRS-C LESC CUSUM MRS-C LESC CUSUM MRS-C LESC
5 2 52.9 ($0.63$) 159.6 ($2.70$) 850.2 ($20.14$) 20.1 ($0.12$) 77.1 ($0.76$) 90.6 ($1.67$) 8.4   ($0.03$) 63.7 ($0.09$) 40.1 ($0.27$)
5 3 37.9 ($0.23$) 122.3 ($1.36$) 722.8 ($23.08$) 14.4 ($0.05$) 76.2 ($0.33$) 85.8 ($1.38$) 6.0   ($0.03$) 57.8 ($0.22$) 42.9 ($1.09$)
10 2 52.5 ($0.33$) 196.3 ($5.59$) 1225.7 ($12.30$) 20.2 ($0.18$) 86.8 ($1.59$) 114.1 ($2.36$) 8.4   ($0.05$) 59.9 ($0.41$) 45.2 ($0.36$)
10 3 38.0 ($0.16$) 153.8 ($2.20$) 1043.9 ($16.91$) 15.8 ($0.10$) 75.0 ($0.44$) 100.2 ($2.76$) 6.0   ($0.03$) 62.7 ($0.30$) 50.3 ($1.04$)
20 2 52.4 ($0.41$) 349.1 ($14.67$) 2072.5 ($43.28$) 20.1 ($0.05$) 106.9 ($0.77$) 185.6 ($8.40$) 8.4   ($0.04$) 63.2 ($0.22$) 47.1 ($0.33$)
20 3 38.1 ($0.28$) 247.5 ($4.05$) 1802.0 ($33.17$) 14.4 ($0.07$) 101.2 ($1.12$) 166.4 ($7.75$) 6.0   ($0.03$) 63.2 ($0.15$) 53.7 ($0.29$)
@@
Table 2 reports the numerical values of the EDD for each detector when its threshold is tuned to yield an ARL of 5000 under the pre-change model. We consider three ambient dimensions $k\in\{5,10,20\}$ and two subspace ranks $d\in\{2,3\}$, across the same three noise variances as in Figure 1. In every configuration, the Exact CUSUM attains the smallest delay, matching its oracle status. MRS-C delivers a delay that is only slightly larger, with the gap narrowing as the signal-to-noise ratio improves (i.e. $\sigma^{2}$ decreases). The Largest-Eigenvalue Shewhart chart again shows substantially higher delays at low SNR, though it becomes more competitive for stronger signals. These results corroborate the trade-offs between ARL and EDD seen in Figure 1 and demonstrate that MRS-C offers a great balance between false-alarm control and rapid detection in multi-rank spiked covariance scenarios.
6.2 Parallel procedure with unknown subspace dimension
We next evaluate the parallel procedure proposed in Section 5 in the setting where the post-change subspace dimension is unknown. In contrast to the experiments in Section 6.1, here multiple MRS-C charts are run in parallel for candidate dimensions, and an alarm is raised when any one of them signals.
We consider the emerging subspace model in ambient dimension $k=20$. Pre-change observations follow $N(0,I_{k})$, and after the change-point fixed at $\tau=500$, post-change observations follow
$x_{t}\sim N(0,I_{k}+U_{d^{\star}}\Lambda_{d^{\star}}U_{d^{\star}}^{T}),\quad\Lambda_{d^{\star}}=I_{d^{\star}},$
where $U_{d^{\star}}\in\mathbb{R}^{k\times d^{\star}}$ has $d^{\star}$ orthonormal columns. We fix the window length at $w=50$, and the candidate dimension set is $\mathcal{D}=\{1,2,\dots,10\}.$ Each individual MRS-C statistic $S_{t}^{(d)}$ is assigned its own threshold $b^{(d)}$, calibrated individually to achieve a per-chart ARL of 50000, ensuring that the overall procedure across ten parallel charts achieves an ARL of 5000 via a Bonferroni-type argument.
The resulting thresholds calibrated by Monte-Carlo simulations and their empirical ARLs are reported in Table 3. Table 3 also summarizes rank-selection frequencies for two post-change scenarios with true ranks $d^{\star}=3$ and $d^{\star}=8$. In both cases, the most frequently selected dimension matches the true post-change dimension, demonstrating that the parallel procedure can infer the latent post-change dimension from the data.
Table 3: Calibrated per-dimension thresholds $b^{(d)}$ achieving ARL $\approx 50000$ for each candidate dimension $d$. The last two rows report empirical rank-selection frequencies from the parallel MRS-C procedure under two different post-change scenarios with true ranks $d^{\star}=3$ and $d^{\star}=8$, respectively. In both cases, the mode of the selected dimension correctly matches the true rank.
$d$ 1 2 3 4 5 6 7 8 9 10
$b^{(d)}$ 40.326 43.594 44.961 46.473 46.731 47.390 48.164 48.220 48.572 48.628
ARL 50254.1 50791.2 49500.2 50096.1 49551.8 49037.4 49983.3 50477.2 49480.2 50332.0
# selected ($d^{\star}=3$) 434 840 876 626 495 407 331 292 234 280
# selected ($d^{\star}=8$) 48 106 177 313 375 484 696 1101 840 664
Beyond identifying the true rank, the parallel strategy must also deliver competitive detection performance. We compare the proposed parallel procedure against the two fixed-rank approaches: MRS-C tuned to the true rank $d^{\star}$ and MRS-C detector with a misspecified dimension that fixes $d=1$. The fixed-rank methods are calibrated to achieve an ARL of 5000.
Detection performance is summarized in Table 4. Across both scenarios, the parallel strategy closely tracks the detector tuned to the true rank and substantially outperforms the misspecified fixed-rank alternative. This demonstrates that uncertainty in the post-change dimension can be handled with little additional cost through the proposed parallel procedure, whereas severe rank misspecification leads to a pronounced loss of sensitivity.
The counts of false alarms before the change-point $\tau$ in Table 4 provide an empirical measure of the effective false-alarm rate. Across both values of $d^{\star}$, the parallel procedure produces substantially fewer premature alarms than the true-rank and misspecified alternatives, reflecting the conservative nature of the Bonferroni calibration. In particular, although all methods are nominally calibrated to $\text{ARL}\approx 5000$, the smaller number of early alarms for the parallel procedure indicates a larger effective ARL in finite samples. This conservatism partially offsets the slight increase in detection delay relative to the true-rank procedure and contributes to the overall robustness of the parallel approach.
Overall, these results demonstrate that the parallel MRS-C procedure provides a robust and practically effective alternative to the true rank known detector, while avoiding the severe penalties associated with rank misspecification. The parallel procedure successfully controls false alarms and reliably detects the change without requiring prior knowledge of the true rank.
Table 4: Comparison of oracle, parallel, and misspecified MRS-C procedures under true ranks $d^{\star}=3$ and $d^{\star}=8$. All methods are calibrated to achieve a global ARL of 5000 under the pre-change distribution. Reported EDDs are computed over runs with $T\geq\tau$ (standard errors in parentheses) .
True rank $d^{\star}$ Dimension selection EDD False alarms before $\tau$
33 True rank ($d=3$) 101.80 ($0.29$) 429
Parallel ($d\in\{1,\dots,10\}$) 107.21 ($0.45$) 185
Misspecified ($d=1$) 128.62 ($0.62$) 399
88 True rank ($d=8$) 68.24 ($0.13$) 448
Parallel ($d\in\{1,\dots,10\}$) 69.99 ($0.12$) 196
Misspecified ($d=1$) 102.15 ($0.43$) 392
6.3 Robotic swarm monitoring
In this section, we qualitatively demonstrate the practical performance of the MRS-C procedure on robotic swarm datasets. We consider both a controlled synthetic swarm-trajectory dataset and a real-world UAV swarm dataset with annotated video frames. These two cases allow us to evaluate the method under complementary conditions, one in which the ground-truth change mechanisms are partially understood, and another in which changes are inferred solely from observed formation transitions.
6.3.1 Synthetic milling data
Figure 2: Snapshot of the synthetic milling swarm configuration at frame $t=6001$. Red dots denote agent positions, and blue arrows indicate instantaneous velocity vectors.
@@ 6.3.1 Synthetic milling data
We consider a synthetic swarm-trajectory dataset comprising $T=10001$ sequential frames of collective behavior, recorded every 0.05s over the physical-time interval 500.00-1000.00s. We index these frames as $t=1,\dots,T$ throughout this section. In each frame, there are $n=100$ agents moving in a 2D plane, and we record each agent’s two-dimensional positions $P_{t}=\{p_{t1},\ldots,p_{tn}\}$ and velocity vectors $V_{t}=\{v_{t1},\ldots,v_{tn}\}$, where $p_{ti},v_{ti}\in\mathbb{R}^{2}$, $i=1,\ldots,n$. Figure 2 shows a representative snapshot of the milling state at time frame $t=6001$. We can construct an affinity graph, where each agent corresponds to a node, and we can observe a time series of nodal features $\hat{X}_{t}=[P_{t};V_{t}]\in\mathbb{R}^{4\times n}$. Let $\tilde{X}_{t}$ be the position-centered and normalized nodal feature values from $\hat{X}_{t}$, which is used as the input for change-point detection algorithms after vectorization.
Figure 3: Detection statistics for the swarm dataset, zoomed in on the physical-time interval 780-820 s, as produced by different change-point detection methods. Both the MRS-C and iso-mirror methods identify the change-point at approximately time 800 s.
@@
We compare MRS-C with two other baseline approaches, including Spectral CUSUM (zhang2023spectral) and the iso-mirror method (chen2023discovering). Figure 3 illustrates the result. MRS-C yields some clearly identifiable change-points, as evidenced by a sharp rise in its detection statistic. Although ground-truth change-time labels are unavailable for this dataset, the timing of the detected change-point closely matches that of Spectral CUSUM and an independent result obtained via the iso-mirror method. Notably, the MRS-C and the iso-mirror method are founded on entirely different statistical and algorithmic principles. We note that the iso-mirror method is an offline procedure that operates on the entire time series retrospectively, whereas MRS-C is designed for real-time sequential detection. The concordance between the detected change-points strongly suggests that both methods capture the same underlying network transition.
6.3.2 UAVSwarm dataset
We evaluated our method on the UAVSwarm dataset (wang2022uavswarm), which contains annotated video sequences of multiple unmanned aerial vehicles (UAVs) operating in swarm configurations under various motion patterns and environmental conditions. In particular, we used the UAVSwarm-13 sequence, which consists of 119 frames. Each frame is provided with ground-truth bounding boxes for all UAVs in view, including their object IDs, image-plane coordinates, and visibility indicators. The number of UAVs in this sequence varies across frames due to occlusions and field-of-view changes. For each frame, we extracted the center coordinates of every agent from the annotated bounding boxes, forming a structured trajectory tensor.
In the UAVSwarm-13 sequence, the UAVs initially maintain a parallel formation, flying in evenly spaced rows with stable relative positions. Around the detected change-point, several agents begin coordinated maneuvers that gradually alter the formation’s geometry. By the mid-transition stage, the swarm reconfigures into a triangular formation, with one UAV leading and others forming staggered trailing positions. Our MRS-C method clearly captures this structural transition, as evidenced by a sharp rise in the MRS-C statistic during the maneuver. Figure 4 shows representative frames and corresponding MRS-C statistics illustrating this process: before the change (parallel), during the transition, and after the new triangular formation is established.
(a) Before change (Frame 3)
(b) Transition onset (Frame 22)
(c) During change (Frame 50)
(d) After change (Frame 79)
Figure 4: Illustration of swarm behavior before, during, and after the detected change-point from the UAVSwarm-13 sequence.
@@ 7 Conclusion
In this paper, we present the Multi-rank Subspace-CUSUM (MRS-C) procedure for the sequential detection of changes in covariance structure in high-dimensional streams. Our method generalizes prior subspace monitoring approaches to handle multiple concurrent spikes, thereby enhancing detection capabilities in scenarios with multiple simultaneous low-rank signals. We established a solid theoretical foundation to guide the selection of key tuning parameters that control the false-alarm rate and minimize detection delay. Analytically, we proved that the proposed detector achieves the same asymptotic order of detection delay as the oracle procedure, with a relative efficiency factor that depends on the uniformity of the signal eigenvalues. We also addressed the practical challenge of unknown subspace dimensions through a parallel monitoring scheme. Numerical experiments on synthetic data and real-world robotic swarm datasets, including UAV formations, confirmed that our method reliably achieves targeted performance and effectively captures complex structural transitions. Future work may extend this framework to non-Gaussian settings or time-varying covariances and explore adaptive strategies for dynamic subspace dimension selection.
@@ Acknowledgement
This work is supported by the Office of Naval Research [N000142412278]. Ira B. Schwartz and Jason Hindes were supported by the Office of Naval Research [N0001425GI101158] and NRL Base funding [N0001424WX00013].
Disclosure of interest
No potential competing interest was reported by the authors.
@@ Appendix A Proofs
For notational simplicity, we here denote by $\hat{u}_{i}$, instead of $\hat{u}_{i,t+w}$, the $i$-th leading unit-norm eigenvector of the sample covariance $\hat{\Sigma}_{t+w}$, which is the $i$th column of $\hat{U}_{t+w}$. Let $\{u_{d+1},\dots,u_{k}\}$ be an arbitrary orthonormal basis of $\mathrm{col}(U)^{\perp}$. Under the spiked covariance model with distinct signal strengths $\{\lambda_{i}\}$, we have the following asymptotic normality of $\hat{u}_{i}$, $i=1,\dots,d$ (anderson1963asymptotic; paul2007asymptotics):
$\sqrt{w}(\hat{u}_{i}-u_{i})\stackrel{{\scriptstyle d}}{{\to}}N(0,\Xi_{i}),$ (A.1)
where the $k\times k$ covariance matrix $\Xi_{i}$ admits the spectral decomposition as
@@
This $\Xi_{i}$ can be rewritten as
$\Xi_{i}=\frac{1+\rho_{i}}{\rho_{i}^{2}}(I_{k}-UU^{T})+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{d}\frac{(1+\rho_{i})(1+\rho_{j})}{(\rho_{i}-\rho_{j})^{2}}u_{j}u_{j}^{T}.$ (A.2)
By the eigenvector CLT (A.1), we may write
$\hat{u}_{i}=u_{i}+\frac{1}{\sqrt{w}}\xi_{i}+o_{p}\!\left(\frac{1}{\sqrt{w}}\right),$
where $\xi_{i}\sim N(0,\Xi_{i})$ is orthogonal with $u_{i}$. Consider the normalization map $g(v)=v/\|v\|$. A second-order Taylor expansion of $g$ around a unit vector $u$ yields
$g(u+h)=u+(I-uu^{\top})h-\frac{1}{2}\big\|(I-uu^{\top})h\big\|^{2}u+o(\|h\|^{2}).$
Taking $h=w^{-1/2}\xi_{i}$ and using $u_{i}^{\top}\xi_{i}=0$, we obtain
$\hat{u}_{i}=u_{i}+\frac{1}{\sqrt{w}}\xi_{i}-\frac{1}{2w}\|\xi_{i}\|^{2}u_{i}+o_{p}\!\left(\frac{1}{w}\right).$
Consequently,
$\hat{u}_{i}\hat{u}_{i}^{\top}=u_{i}u_{i}^{\top}+\frac{1}{\sqrt{w}}\big(u_{i}\xi_{i}^{\top}+\xi_{i}u_{i}^{\top}\big)+\frac{1}{w}\!\left(\xi_{i}\xi_{i}^{\top}-\|\xi_{i}\|^{2}u_{i}u_{i}^{\top}\right)+o_{p}\!\left(\frac{1}{w}\right),$
and hence
$\mathbb{E}_{0}(\hat{u}_{i}\hat{u}_{i}^{\top})=u_{i}u_{i}^{\top}+\frac{1}{w}\!\left(\Xi_{i}-\operatorname{tr}(\Xi_{i})\,u_{i}u_{i}^{\top}\right)+o\!\left(\frac{1}{w}\right).$ (A.3)
From (A.2), we have
$\mbox{tr}(\Xi_{i})=(k-d)\frac{1+\rho_{i}}{\rho_{i}^{2}}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{d}\frac{(1+\rho_{i})(1+\rho_{j})}{(\rho_{i}-\rho_{j})^{2}}.$ (A.4)
Recall that $\hat{u}_{i}$ is independent with $x_{t}$. Using (A.3) and (A.4), we get